OpenStudy (anonymous):

solve the equation x/3-x/7=2

OpenStudy (anonymous):

When you are solving (there is an = sign) you find the common denominator but instead of putting it on the bottom you multiply "EVERY" term by that number. So you common denominator is 3 * 7 = 21 so multiply each term by 21 21(x/3) - 21(x/7) = 21 * 2 The 3 cancels up into the 21, the 7 up into the 21 7x - 3x = 42 4x = 42 x = 10.5

OpenStudy (anonymous):

:)